# Improving sampling of binding free energy differences between covalently bound ligands in alternate binding pockets using MT-REXEE

### v1

@@ 4.3 Configurational and Alchemical Free Energy
This sampling of both pockets during the same simulation allows us to achieve two goals: Estimating the relative free energy as a function of chain length independently in each pocket and estimating the configurational free energy difference for a given chain length substrate in pocket A versus pocket B. The solvent reference state for all alchemical binding free energies is the acyl chain attached to ACP in solvent as described in the Methods section (Figure 12).
We first ensured that the overlap between adjacent $\lambda$ states was sufficient for all transformations both in solvent (Figure 13) and in complex (Figure 14). We used the same $\lambda$ state spacing for both the EE and MT-REXEE simulations, as well as the same initial configurations for all replicas.
@@
The EE simulations were initiated with the same chain conformations as the MT-REXEE simulations in pocket A. Since the EE transformations initialized in pocket A are only able to sample pocket A, we can only make comparisons to the alchemical free energy transformations in pocket A. We observe a significant difference between the free energy estimate for the C4 to C6, C6 to C8, and C12 to C14 transformations when using MT-REXEE compared to EE (Figure 4A). It is not unexpected that some transformations would produce significantly different FE estimates as even within a given pocket the configurational space sampled with each of the two methods is expected to vary greatly. This is less likely to affect the known flexible transformations like C2 to C4, which would be expected to readily sample the conformational space in both methods. The degree of sampling with a given binding pocket is also less likely to affect the longest substrates which have fewer accessible conformations which fit the full acyl chain. Overall, these results are consistent with the significantly different conformational sampling expected with MT-REXEE compared to EE. We would worry about MT-REXEE introducing additional error into the calculations producing these differences; however, the fact that the C2 to C4 transformation, the one transformation that we would expect EE alone to sample effectively, is statistically identical between the two methods suggests no errors are introduced.
@@
Figure 4: The convergence of binding free energy simulations initiated in pockets A and B demonstrates MT-REXEE can increase sampling of the conformational space. (A) Comparison of estimates from individual independent EE calculations with estimates of the same transformations in MT-REXEE simulations. MT-REXEE’s significantly improved sampling leads to moderate differences between the two methods, with MT-REXEE expected to be more accurate. Estimates for binding free energy using MT-REXEE with the initial acyl chain configurations in either pocket A (brown) or pocket B (orange) for the calculation of the RFE in pocket A (B) and pocket B (C), showing independence of starting configuration. (D) Cumulative relative binding free energy can be obtained for all chain lengths relative to a C2 chain for both pockets using simulations initiated in both pockets A and B (Figure 15). Thus, from a single set of MT-REXEE simulations, we can obtain the full chain length and pocket dependence on binding FE. For all figures, we plot the mean of n=3 (n=6 for D) replicas and error bars represent the standard error of the mean.
@@
We also performed MT-REXEE simulations initiated in pocket B to assess the convergence of the free energy estimates between the pockets, when the simulations were started in pocket A versus started in pocket B. After 100 ns of MT-REXEE simulation, the estimates were within statistical error regardless of whether they started in pocket A (Figure 4B) or pocket B (Figure 4C) to within statistical error, with the possible exception of the C14 to C16 transformation initialized in pocket A. A single simulation of this transformation which was initiated in pocket B experienced significantly less sampling in pocket A than the other two replicas. This contributes significantly to this deviation. As the simulation get longer, additional metastable states may emerge, so increase in sampling difficulty for very long chains is not unexpected.
@@
The G107M variant is known to prevent the synthesis of fatty acids past 8 carbons in length in FabF, an enzyme isoform which does not have pocket B. The available experimental data on activity, along with the location of the mutations within the binding pocket, suggest that binding to pocket A is likely to be destabilized for substrates longer than 8 carbons in length (Figure 5A). We applied the MT-REXEE approach to measure binding free energy estimates for all chain lengths in both pocket A and pocket B to determine if this mutation affects the stability of the covalent intermediate state for chain lengths greater than 8 and to quantify that effect.
We compared the relative binding free energy values for all chain lengths in pocket A and pocket B between the WT and G107M FabB variants (Figure 5B–C). We observed that the G107M mutation appears to have a stabilizing affect on the complex for C2–6 chain lengths, as shown by the negative free energy difference between the mutations, shown in Figure 5D. The larger residue at position 107 is able to form additional non-bonded interactions within the pocket as well as with the short acyl chains which may explain this stabilizing affect. However, at the C6 to C8 transformation we see a significant destabilization of pocket A in the G107M variant which is not present for the WT as shown by the positive free energy difference in figure 5D. This difference is particularly evident for pocket A but not pocket B as shown by the difference in relative binding free energy between pocket A and pocket B ($\Delta G_{A}-\Delta G_{B}$). In the WT variant, chain growth into pocket A is more stable than pocket B until 6 carbons (subfigure D), the two pockets are statistically identical until 10 carbons, and then for 12 carbons or greater pocket B is more stable for each chain growth step into the pocket. However, in the G107M variant, pocket B is significantly more stable while growing the chain from 6 to 10 carbons (Figure 5D). This stark difference in relative binding free energies between pockets A and B coincides with the chain length at which binding in pocket A is predicted to be destabilized based on experimental evidence.
The focus of this work is on establishing the utility of the MT-REXEE method in increasing conformational sampling and predicting physically realistic complex conformations, but briefly we will review how restriction of loop motions provides a physical explanation for the observed differences in FE profile for the G107M FabB variant (Figure 16). Structural analyses of the trajectories suggest that the reason for this destabilization is that the G107M mutation restricts the accessible binding pocket for substrates longer than 8 carbons.
@@
There is a loop consisting of residues 105–110 in FabB which has two conformations. This loop reorientation occurs only for substrate C8 or longer in the WT FabB simulations. The G107M mutations prevent loop reorientation, which restricts the conformational space accessible by the acyl chain. Additionally, only in the reoriented conformation is this loop able to form stable non-bonded interactions with acyl chain carbons 7–9 which could be responsible for the destabilization of growing in C8 and C10 acyl substrates to FabB in the G107M variant. Once the chain has grown to 10 carbons we no longer see this additional cost for further extending the chain suggesting that the penalty is specifically for growing C8 and C10 chains which would not longer be included for growing in C12 or larger acyl chains. Further experimentation is necessary to validate this mechanism as being the primary contribution to this thermodynamic effect, but we have demonstrated the ability of the MT-REXEE to provide both thermodynamic quantification as well as a physical explanation for experimentally observed protein behavior in highly flexible systems.
@@ 5 Discussion and Conclusions
The MT-REXEE method increases conformational sampling for sets of alchemical transformations in which prohibitively high free energy barriers are present at only a subset of the states of interest. We use the FabB—acyl complex to demonstrate the potential of MT-REXEE to increase conformational sampling. In the FabB—acyl complex, two distinct pockets can be occupied by the acyl chain. Pocket B was inaccessible using standard MD simulations initiated in pocket A for any chain length longer than 4 carbons. MT-REXEE allowed for significantly enhanced sampling between the two pockets within a single trajectory for each transformation. This enhanced sampling has then been leveraged to quantify the difference in binding free energy surfaces between pocket A and pocket B for the G107M FabB variant, which has a known effect of preventing medium chain synthesis through pocket A.
@@
This work also highlights the new exchange method, random range (RR), in which conformational exchanges are allowed at any frame within a given percentage of the trajectory rather than the last frame of an iteration. This improvement removes the necessity of including redundant or up-weighted end states which was previously necessary. We verified that this novel exchange method does not introduce additional error into the calculations through comparison of binding free energy estimates to MUP1 using both EE and MT-REXEE using the previous exhaustive method. All three methods produce statistically equivalent binding FE estimates, but we observe significantly faster convergence using MT-REXEE with the RR method. The RR procedure significantly reduces the round trip times in alchemical space while also increasing relative sampling of intermediate $\lambda$ states with the same trajectory length. Decreasing the round trip time increases the efficiency of the MT-REXEE method, which is of growing importance when simulating larger systems like that of the FabB–acyl complex. There is a trade-off when using the RR method in which an average of $\sim$25% of each iteration’s sampling (for the settings used in this study) is not included in the trajectories leading to conformational swaps. This sampling is still included in FE and trajectory analysis and is thus not lost, but will reduce the degree of decorrelation of the pre-swap configurations because of decreased total sampling time along continuous trajectories.
@@
For the FabB–acyl complex, we simulate the covalent intermediate state. This state was selected to simplify the test system, compared to the FabB–acyl-ACP complex. Much like the formation of covalent bonds in intermediate states of enzymatic processes, covalent drugs initiate the formation of covalent bonds with their protein targets. Alchemical free energy estimates for reversible covalent drug compounds have been successfully computed, but often for relatively rigid ligands 4. The conformational sampling problem caused by increased conformational flexibility in the ligand is exacerbated by the anchor provided by the covalent bond which limits potential ligand and protein target rearrangements. The application of MT-REXEE to the FabB–acyl system demonstrates that converged estimates for the binding free energy for flexible covalent ligands can be achieved with reasonable computational resources without needing to identify or define collective variables.
@@
The Gibbs FE differences are not directly additive $\Delta G_{ij}\neq\Delta G_{ij,A}+\Delta G_{ij,B}$ as this would assume equal contributions from both configurational states. To get a direct expression for the free energy difference between states $i$ and $j$ regardless of configurational state ($\Delta G_{ij}$) in terms of $\Delta G_{ij,A}$ and $\Delta G_{ij,B}$ we start by rearranging equation 2 to give us equation 9 and 10. For simplicity’s sake, we utilize the variables $A_{i}$, $A_{j}$, $B_{i}$, and $B_{j}$ as defined in equations 9 and 10.
@@
With this reduced representation we can rewrite equation 1 as $\Delta G_{ij}=-k_{B}T\ln{\frac{A_{i}+B_{i}}{A_{j}+B_{j}}}$. When we substitute in the values for Equations 9 and 10 we get
@@
Figure 6: This diagram shows how we can discretize trajectories based on whether they are sampling from conformational space A(orange) or B(pink). Over the course of a single iteration the system will perform Monte Carlo transitions between $\lambda$ states which is represented by movement along the y-axis. The system will also freely move in configurational space between space A and B shown by changes in vertical line color. In analysis we can separate the trajectory into frames which sample in configurational space A and B (dashed red lines). We concatenate the energies for all $\lambda$ states sampled in configurational space A and B and then perform separate MBAR analysis to obtain the free energy estimates.
@@
Figure 7: The separation of the free energy into conformational and alchemical free energy produces three distinct free energy cycles. One cycle follows the blue and pink arrows computing the relative free energy between Cn to Cn+2. The other two cycles follow the pink and either light or dark orange arrows computing the relative free energy for Cn between pockets A and B and separately Cn+2 and pockets A or B.
@@
Figure 8: (A) We highlight the location of our reference residue THR300 (blue) position relative to the chain conformation in pocket A (orange) compared to pocket B (purple). (B) We plot the distance distributions in a set of trajectories with chain lengths ranging from 2 to 16 carbons and known to be in pocket A or B. The independent peaks demonstrate that this is an ideal metric for determining quantitatively defining the two pockets.
@@
Figure 9: We compare the Wang-Landau weight equilibration times for all MUP1 small molecule ligand transformations performed both in solvent (A) and in the protein complex (B). The weight equilibration times for the MT-REXEE method with the Random Range equilibration method and no redundant end-states results in either statistically equivalent or lower equilibration times.
@@
Figure 10: Here we use the absolute value of the difference in FE between using the first and last 50% of the trajectory following Wang-Landau weight equilibration. If the estimates are closer together at 50% we can conclude that the free energy estimates are converging more quickly which is likely a result of increased conformational sampling. (A) In the solvent simulations with the MUP1 ligands we see no real significant difference in the rate of convergence between methods. However in the protein complex simulations (B), we see a significant decrease in the difference between the forward and reverse estimates which indicates faster conformational sampling with MT-REXEE. (C) We also use the standard error between replicates as a measure of convergence for the free energy simulations which shows a significant decrease for the exhaustive method and an even more drastic decrease for MT-REXEE RR method.
@@
Figure 11: (A) The round trip time is the time it takes for a conformation to swap through all parallel simulations and return to the original $\lambda$ state. The round trip times are divided into the first and last half of the trajectories. We see a significant increase in the round trip time in the latter half of the simulation. (B) We compare the $\lambda$ ratio for all $\lambda$ states in each transformation in the first and second half of the trajectory. This is computed as $R=\frac{N_{i}}{N_{avg}}$ where $N_{i}$ is the number of frames sampled at a given $\lambda$ states, and $N_{avg}$ is the average frames sampled for all $\lambda$ states. In an ideal situation R = 1 because every state is sampled evenly. A value of R¡1 means the state is under sampled and R¿1 means the state is over-sampled. We observe a corresponding shift in the observed $\lambda$ ratio in the latter half of the simulation. This is possibly because the sampled conformational space is deviating significantly from the space sampled as the weights were equilibrating.
@@
Figure 12: As the solvent reference state here we utilize the ACP with a sequestered acyl chain. We show centroid structures from the MT-REXEE simulations of the solvent state for C2-4 (A) and the C14-16 (B) states.
@@
Figure 13: We measure the configurational overlap between $\lambda$ states in the ACP solvent simulations for transformations (A) A-B, (B) B-C, (C) C-D, (D) D-E, (E) E-F, (F) F-G, (G) G-H. We optimized the $\lambda$ state spacing in order to ensure off-diagonal values of at least 0.05.
@@
Figure 14: We measure the configurational overlap between $\lambda$ states in the FabB complex simulations for transformations (A) A-B, (B) B-C, (C) C-D, (D) D-E, (E) E-F, (F) F-G, (G) G-H. We optimized the $\lambda$ state spacing in order to ensure off-diagonal values of at least 0.05.
@@
Figure 15: We present an example schematic for computing the cumulative FE difference to a single reference point as used in figure 4. In this example we compute the relative free energy difference between a C2 chain in pocket A and a C6 chain in pocket B. We add the configurational FE difference between pockets A and B for a C2 chain ($\Delta G_{conf,0}$) to the alchemical free energy difference of performing the C2 to C4, C4 to C6, and C6 to C8 transformations in pocket B ($\Delta G_{0,B}$, $\Delta G_{1,B}$, $\Delta G_{2,B}$ respectively).
@@
Figure 16: The G107M mutations is observed to add a free energy cost for growing the acyl chain into pocket A for C6-8 and C8-10 transformations using MT-REXEE simulations. One physical explanation is that the restriction of conformational space means the blue C12 conformation is accessible only in the WT while the C12 conformation in pink is the primary accessible conformation in the G107M FabB variant. This chain rearrangement coincides with the movement of a loop consisting of residues 105-110. This loop rearrangement is only possible in the WT and not the G107M variant and these loop residues form non-bonded interactions with carbons 7–10 on the acyl chain only in the WT conformation.

### v2

@@ 4.3 Configurational and Alchemical Free Energy
This sampling of both pockets during the same simulation allows us to achieve two goals: Estimating the relative free energy as a function of chain length independently in each pocket and estimating the configurational free energy difference for a given chain length substrate in pocket A versus pocket B. The solvent reference state for all alchemical binding free energies is the acyl chain attached to ACP in solvent as described in the Methods section (Figure S7).
@@
The EE simulations were initiated with the same chain conformations as the MT-REXEE simulations in pocket A. Since the EE transformations initialized in pocket A are only able to sample pocket A, we can only make comparisons to the alchemical free energy transformations in pocket A. We observe a significant difference between the free energy estimate for the C4 to C6, C6 to C8, and C12 to C14 transformations when using MT-REXEE compared to EE (Figure 4A). It is not unexpected that some transformations would produce significantly different FE estimates as even within a given pocket the configurational space sampled with each of the two methods is expected to vary greatly. This is less likely to affect the known flexible transformations like C2 to C4, which would be expected to readily sample the conformational space in both methods. The degree of sampling with a given binding pocket is also less likely to affect the longest substrates which have fewer accessible conformations that fit the full acyl chain. Overall, these results are consistent with the significantly different conformational sampling expected with MT-REXEE compared to EE. We would worry about MT-REXEE introducing additional error into the calculations producing these differences; however, the fact that the C2 to C4 transformation, the one transformation that we would expect EE alone to sample effectively, is statistically identical between the two methods suggests no errors are introduced.
@@
Figure 4: The convergence of binding free energy simulations initiated in pockets A and B demonstrates MT-REXEE can increase sampling of the conformational space. (A) Comparison of estimates from individual independent EE calculations with estimates of the same transformations in MT-REXEE simulations. MT-REXEE’s significantly improved sampling leads to moderate differences between the two methods, with MT-REXEE expected to be more accurate. Estimates for binding free energy using MT-REXEE with the initial acyl chain configurations in either pocket A (brown) or pocket B (orange) for the calculation of the RFE in pocket A (B) and pocket B (C), showing independence of starting configuration. (D) Cumulative relative binding free energy can be obtained for all chain lengths relative to a C2 chain for both pockets using simulations initiated in both pockets A and B (Figure S10). Thus, from a single set of MT-REXEE simulations, we can obtain the full chain length and pocket dependence on binding FE. For all figures, we plot the mean of n=3 (n=6 for D) replicas and error bars represent the standard error of the mean.
@@
We also performed MT-REXEE simulations initiated in pocket B to assess the convergence of the free energy estimates between the pockets, when the simulations were started in pocket A versus started in pocket B. After 100 ns of MT-REXEE simulation, the estimates were within statistical error regardless of whether they started in pocket A (Figure 4B) or pocket B (Figure 4C), with the possible exception of the C14 to C16 transformation initialized in pocket A. A single simulation of this transformation, when initiated in pocket B, showed significantly less sampling in pocket A than the other two replicas, and contributes significantly to this deviation. If simulations were extended additional metastable states may emerge, so greater sampling difficulty for very long chains is expected.
@@
In E. coli, FabB and FabF are homologs that show different responses to mutations. In FabF, I108M prevents the synthesis of acyl chains longer than 8 carbons; in FabB, G107M (the analogous mutation) is more permissive but still shifts product profiles toward shorter chains 24. These effects suggest that only FabB can access pocket B, which permits binding of longer chains when pocket A is obstructed. (Figure 5A). Focusing on FabB, we applied the MT-REXEE approach to measure binding free energy estimates for all chain lengths in pockets A and B to determine how G107M affects the stability of the covalent intermediate for different chain lengths (Figure 5B–C). Interestingly, the G107M mutation appears to stabilize the binding of C2–6 chain lengths, as shown by the negative free energy difference between the mutations, shown in Figure 5D. The larger residue at position 107 forms additional non-bonded interactions within the pocket as well as with the short acyl chains, which may explain this stabilizing affect. For the C6 to C8 transformation, however, we see a significant destabilization of pocket A in the G107M variant, relative to WT, as shown by the positive free energy difference 5D. This difference is particularly evident for pocket A but not for pocket B, as shown by the difference in relative binding free energy between pockets A and B ($\Delta G_{A}-\Delta G_{B}$). In the WT variant, chain growth into pocket A is more stable than pocket B until 6 carbons, the two pockets are statistically identical until 10 carbons, and then for 12 carbons or greater, pocket B is more stable for each chain growth step. However, for G107M, pocket B is significantly more stable while growing the chain from 6 to 10 carbons (Figure 5D). This stark difference in relative binding free energies between pockets A and B coincides with the chain length at which binding in pocket A is predicted to be destabilized based on experimental evidence.
@@ 5 Discussion and Conclusions
The MT-REXEE method increases conformational sampling for sets of alchemical transformations in which prohibitively high free energy barriers are present at only a subset of the states of interest. In this paper, we used the FabB-acyl complex to demonstrate the potential of MT-REXEE to increase conformational sampling. In the FabB-acyl complex, the acyl chain could access two distinct pockets, consistent with previous crystallographic analyses. Pocket B was inaccessible using standard MD simulations initiated in pocket A for any chain length longer than 4 carbons. MT-REXEE allowed for significantly enhanced sampling between the two pockets within a single trajectory for each transformation. This enhanced sampling enabled quantification of the difference in binding free energy surfaces between pocket A and pocket B for the G107M FabB variant, which shifts product profiles toward short chains but still permits medium chain binding.
@@
This work also highlights the new exchange method, random range (RR), in which conformational exchanges are allowed at any frame within a given percentage of the trajectory rather than the last frame of an iteration. This improvement removes the necessity of including redundant or up-weighted end states which was previously necessary. We verified that this novel exchange method does not introduce additional error into the calculations through comparison of binding free energy estimates to MUP1 using both EE and MT-REXEE using the previous exhaustive method. All three methods produce statistically equivalent binding FE estimates, but we observe significantly faster convergence using MT-REXEE with the RR method. The RR procedure significantly reduces the round-trip times in alchemical space while also increasing relative sampling of intermediate $\lambda$ states with the same trajectory length. Decreasing the round-trip time increases the efficiency of the MT-REXEE method, which is of growing importance when simulating larger systems like that of the FabB–acyl complex. There is a trade-off when using the RR method in which an average of $\sim$25% of each iteration’s sampling (for the settings used in this study) is not included in the trajectories leading to conformational swaps. This sampling is still included in FE and trajectory analysis and is thus not lost, but will reduce the degree of decorrelation of the pre-swap configurations because of decreased total sampling time along continuous trajectories.
@@
For the FabB-acyl complex, we simulated the covalent intermediate state. This state was selected to simplify the test system, compared to the FabB–acyl-ACP complex. Much like the formation of covalent bonds in intermediate states of enzymatic processes, covalent drugs initiate the formation of covalent bonds with their protein targets. Alchemical free energy estimates for reversible covalent drug compounds have been successfully computed, but often for relatively rigid ligands 30. The conformational sampling problem caused by increased conformational flexibility in the ligand is exacerbated by the anchor provided by the covalent bond which limits potential ligand and protein target rearrangements.
This large highly flexible system does pose important unsolved challenges for the MT-REXEE method. The Wang-Landau weight equilibration times can exceed 100 ns thus requiring significant additional computational resources due to the complex nature of the FE landscape. This problem could be mitigated by implementing educated initial guesses for weights and using adaptive sampling methods like SAMS 39. Additionally, handling multiple simulations that need to be exchanged on the order of minutes, although removing complexity from simulation code, still requires a complex workflow that is not optimally managed by sub-process calls in Python. Workflow software improvements for robustly managing large numbers of biomolecular simulations at scale would allow for simulations to be run much more easily 3. The current framework is functionally limited to running on a single CPU node and is thus hardware-limited.
Despite these limitations, MT-REXEE provides a unique methodology for completely unbiased sampling of alternate conformational states. The intended application for large-flexible systems is not high throughput analysis but rather in-depth analysis of challenging systems for which other methods are either insufficient or require assumptions for which sufficient data is not available to support. This was the case for the FabB-acyl complex system and MT-REXEE simulations were able to provide valuable insights on occupancy of alternate pockets as a mechanism for mutation resistance.
@@
The Gibbs FE differences are not directly additive $\Delta G_{ij}\neq\Delta G_{ij,A}+\Delta G_{ij,B}$ as this would assume equal contributions from both configurational states. To get a direct expression for the free energy difference between states $i$ and $j$ regardless of configurational state ($\Delta G_{ij}$) in terms of $\Delta G_{ij,A}$ and $\Delta G_{ij,B}$ we start by rearranging equation 2 to give us equation 7 and 8. For simplicity’s sake, we utilize the variables $A_{i}$, $A_{j}$, $B_{i}$, and $B_{j}$ as defined in equations 7 and 8.
@@
With this reduced representation we can rewrite equation 1 as $\Delta G_{ij}=-k_{B}T\ln{\frac{A_{i}+B_{i}}{A_{j}+B_{j}}}$. When we substitute in the values for Equations 7 and 8 we get
@@
Figure S1: This diagram shows how we can discretize trajectories based on whether they are sampling from conformational space A(orange) or B(pink). Over the course of a single iteration the system will perform Monte Carlo transitions between $\lambda$ states which is represented by movement along the y-axis. The system will also freely move in configurational space between space A and B shown by changes in vertical line color. In analysis we can separate the trajectory into frames which sample in configurational space A and B (dashed red lines). We concatenate the energies for all $\lambda$ states sampled in configurational space A and B and then perform separate MBAR analysis to obtain the free energy estimates.
@@
Figure S2: The separation of the free energy into conformational and alchemical free energy produces three distinct free energy cycles. One cycle follows the blue and pink arrows computing the relative free energy between Cn to Cn+2. The other two cycles follow the pink and either light or dark orange arrows computing the relative free energy for Cn between pockets A and B and separately Cn+2 and pockets A or B.
@@
Figure S3: (A) We highlight the location of our reference residue THR300 (blue) position relative to the chain conformation in pocket A (orange) compared to pocket B (purple). (B) We plot the distance distributions in a set of trajectories with chain lengths ranging from 2 to 16 carbons and known to be in pocket A or B. The independent peaks demonstrate that this is an ideal metric for determining quantitatively defining the two pockets.
@@
Figure S4: We compare the Wang-Landau weight equilibration times for all MUP1 small molecule ligand transformations performed both in solvent (A) and in the protein complex (B). The weight equilibration times for the MT-REXEE method with the Random Range equilibration method and no redundant end-states results in either statistically equivalent or lower equilibration times.
@@
Figure S5: Here we use the absolute value of the difference in FE between using the first and last 50% of the trajectory following Wang-Landau weight equilibration. If the estimates are closer together at 50% we can conclude that the free energy estimates are converging more quickly which is likely a result of increased conformational sampling. (A) In the solvent simulations with the MUP1 ligands we see no real significant difference in the rate of convergence between methods. However in the protein complex simulations (B), we see a significant decrease in the difference between the forward and reverse estimates which indicates faster conformational sampling with MT-REXEE. (C) We also use the standard error between replicates as a measure of convergence for the free energy simulations which shows a significant decrease for the exhaustive method and an even more drastic decrease for MT-REXEE RR method.
@@
Figure S6: (A) The round trip time is the time it takes for a conformation to swap through all parallel simulations and return to the original $\lambda$ state. The round trip times are divided into the first and last half of the trajectories. We see a significant increase in the round trip time in the latter half of the simulation. (B) We compare the $\lambda$ ratio for all $\lambda$ states in each transformation in the first and second half of the trajectory. This is computed as $R=\frac{N_{i}}{N_{avg}}$ where $N_{i}$ is the number of frames sampled at a given $\lambda$ states, and $N_{avg}$ is the average frames sampled for all $\lambda$ states. In an ideal situation R = 1 because every state is sampled evenly. A value of R¡1 means the state is under sampled and R¿1 means the state is over-sampled. We observe a corresponding shift in the observed $\lambda$ ratio in the latter half of the simulation. This is possibly because the sampled conformational space is deviating significantly from the space sampled as the weights were equilibrating.
@@
Figure S7: As the solvent reference state here we utilize the ACP with a sequestered acyl chain. We show centroid structures from the MT-REXEE simulations of the solvent state for C2-4 (A) and the C14-16 (B) states.
@@
Figure S8: We measure the configurational overlap between $\lambda$ states in the ACP solvent simulations for transformations (A) A-B, (B) B-C, (C) C-D, (D) D-E, (E) E-F, (F) F-G, (G) G-H. We optimized the $\lambda$ state spacing in order to ensure off-diagonal values of at least 0.05.
@@
Figure S9: We measure the configurational overlap between $\lambda$ states in the FabB complex simulations for transformations (A) A-B, (B) B-C, (C) C-D, (D) D-E, (E) E-F, (F) F-G, (G) G-H. We optimized the $\lambda$ state spacing in order to ensure off-diagonal values of at least 0.05.
@@
Figure S10: We present an example schematic for computing the cumulative FE difference to a single reference point as used in figure 4. In this example we compute the relative free energy difference between a C2 chain in pocket A and a C6 chain in pocket B. We add the configurational FE difference between pockets A and B for a C2 chain ($\Delta G_{conf,0}$) to the alchemical free energy difference of performing the C2 to C4, C4 to C6, and C6 to C8 transformations in pocket B ($\Delta G_{0,B}$, $\Delta G_{1,B}$, $\Delta G_{2,B}$ respectively).
@@
Figure S11: The G107M mutations is observed to add a free energy cost for growing the acyl chain into pocket A for C6-8 and C8-10 transformations using MT-REXEE simulations. One physical explanation is that the restriction of conformational space means the blue C12 conformation is accessible only in the WT while the C12 conformation in pink is the primary accessible conformation in the G107M FabB variant. This chain rearrangement coincides with the movement of a loop consisting of residues 105-110. This loop rearrangement is only possible in the WT and not the G107M variant and these loop residues form non-bonded interactions with carbons 7–10 on the acyl chain only in the WT conformation.